The following information pertains to Austin, Inc. and Huston Company:

Account Title Austin Huston
Current assets $ 40,000 $ 62,500
Total assets 480,000 480,000
Current liabilities 12,000 31,250
Total liabilities 290,000 376,000
Stockholders’ equity 190,000 104,000
Interest expense 15,800 18,400
Income tax expense 31,600 28,800
Net income 70,000 70,700


Required
a-1. Compute each company’s debt-to-assets ratio, current ratio, and times interest earned (EBIT must be computed). (Round your "Debt-to-assets ratio" and "Times interest earned" to 1 decimal place and current ratio answers to 2 decimal places.)

Compute each company’s return-on-equity ratio and return-on-assets ratio. Use EBIT instead of net income when computing the return-on-assets ratio. (Round your answers to 2 decimal places.)
Answer #1

a) Calculate following :

Formula Austin Huston
Debt to assets ratio Total liabilities/Total assets 290000/480000 = 0.6 Times 376000/480000 = 0.8
Current ratio Current assets/current liabilities 40000/12000 = 3.33 Times 62500/31250 = 2.00 Times
Times interest earned EBIT/Interest expense (70000+15800+31600)/15800 = 7.4 Times (70700+18400+28800)/18400 = 6.4 Times

b) Calculate following

Austin company Huston company
Return on equity Net income/Equity 70000/190000 = 36.84% 70700/104000 = 67.98%
Return on assets EBIT/Total assets 117400/480000 = 24.46% 117900/480000 = 24.56%